# Complexity Analysis of Normalizing Constant Estimation: from Jarzynski Equality to Annealed Importance Sampling and beyond

### v1

@@ Abstract
Given an unnormalized probability density $\pi\propto\mathrm{e}^{-V}$, estimating its normalizing constant $Z=\int_{\mathbb{R}^{d}}\mathrm{e}^{-V(x)}\mathrm{d}x$ or free energy $F=-\log Z$ is a crucial problem in Bayesian statistics, statistical mechanics, and machine learning. It is challenging especially in high dimensions or when $\pi$ is multimodal. To mitigate the high variance of conventional importance sampling estimators, annealing-based methods such as Jarzynski equality and annealed importance sampling are commonly adopted, yet their quantitative complexity guarantees remain largely unexplored. We take a first step toward a non-asymptotic analysis of annealed importance sampling. In particular, we derive an oracle complexity of $\widetilde{O}\left(\frac{d\beta^{2}{\mathcal{A}}^{2}}{\varepsilon^{4}}\right)$ for estimating $Z$ within $\varepsilon$ relative error with high probability, where $\beta$ is the smoothness of $V$ and $\mathcal{A}$ denotes the action of a curve of probability measures interpolating $\pi$ and a tractable reference distribution. Our analysis, leveraging Girsanov theorem and optimal transport, does not explicitly require isoperimetric assumptions on the target distribution. Finally, to tackle the large action of the widely used geometric interpolation of probability distributions, we propose a new normalizing constant estimation algorithm based on reverse diffusion samplers and establish a framework for analyzing its complexity.
Keywords: Normalizing constant, free energy, Jarzynski equality, annealed importance sampling, reverse diffusion samplers.
@@ 1 Introduction
We study the problem of estimating the normalizing constant $Z=\int_{\mathbb{R}^{d}}\mathrm{e}^{-V(x)}\mathrm{d}x$ of an unnormalized probability density function (p.d.f.) $\pi\propto\mathrm{e}^{-V}$ on $\mathbb{R}^{d}$, so that $\pi(x)=\frac{1}{Z}\mathrm{e}^{-V(x)}$. The normalizing constant appears in various fields: in Bayesian statistics, when $\mathrm{e}^{-V}$ is the product of likelihood and prior, $Z$ is also referred to as the marginal likelihood or evidence (Gelman et al., 2013); in statistical mechanics, when $V$ is the Hamiltonian [^1] Up to a multiplicative constant $\beta=\frac{1}{k_{\mathrm{B}}T}$ known as the thermodynamic beta, where $k_{\mathrm{B}}$ is the Boltzmann constant and $T$ is the temperature. When borrowing terminologies from physics, we ignore this quantity for simplicity., $Z$ is known as the partition function, and $F:=-\log Z$ is called the free energy (Chipot and Pohorille, 2007, Lelièvre et al., 2010, Pohorille et al., 2010). The task of normalizing constant estimation has numerous applications, including computing log-likelihoods in probabilistic models (Sohl-Dickstein and Culpepper, 2012), estimating free energy differences (Lelièvre et al., 2010), and training energy-based models in generative modeling (Song and Kingma, 2021, Carbone et al., 2023, Sander et al., 2025). It is challenging in high dimensions or when $\pi$ is multimodal (i.e., $V$ has a complex landscape).
@@
Conventional approaches based on importance sampling (Meng and Wong, 1996) are widely adopted to tackle this problem, but they suffer from high variance due to the mismatch between target and proposal distributions when the target distribution is complicated (Chatterjee and Diaconis, 2018). To alleviate this issue, the technique of annealing tries constructing a sequence of intermediate distributions that bridge these two distributions, which motivates several popular methods including path sampling (Chen and Shao, 1997, Gelman and Meng, 1998), annealed importance sampling (AIS, Neal (2001)), and sequential Monte Carlo (SMC, Doucet et al. (2000), Del Moral et al. (2006), Syed et al. (2024)) in statistics literature, as well as thermodynamic integration (TI, Kirkwood (1935)) and Jarzynski equality (JE, Jarzynski (1997), Ge and Jiang (2008), Hartmann et al. (2019)) in statistical mechanics literature. In particular, JE points out the connection between the free energy difference between two states and the work done over a series of trajectories linking these two states, while AIS constructs a sequence of intermediate distributions and estimates the normalizing constant by importance sampling over these distributions. These two methods are our primary focus in this paper.
@@
Despite the empirical success of annealing-based methods (Ma et al., 2013, Krause et al., 2020, Mazzanti and Romero, 2020, Yasuda and Takahashi, 2022, Chen and Ying, 2024, Schönle et al., 2024), the theoretical understanding of their performance is still limited. Existing works for importance sampling mainly focus on the asymptotic bias and variance of the estimator (Meng and Wong, 1996, Gelman and Meng, 1998), while works on JE usually simplify the problem by assuming the work follows simple distributions (e.g., Gaussian or gamma) (Echeverria and Amzel, 2012, Arrar et al., 2019). Moreover, only analyses asymptotic in the number of particles derived from central limit theorem exist (Lelièvre et al., 2010, Sec. 4.1). In this paper, we aim to establish a rigorous non-asymptotic analysis of estimators based on JE and AIS, while introducing minimal assumptions on the target distribution. Moreover, we also propose a new algorithm based on reverse diffusion samplers to tackle a potential shortcoming of AIS.
Contributions.
Our key technical contributions are summarized as follows.
  • We discover a novel strategy for analyzing the complexity of normalizing constant estimation, applicable to a wide range of target distributions (see Assumps. 2 and 1) that may not satisfy isoperimetric conditions such as log-concavity.
  • In Sec. 4, we study JE and prove an upper bound on the time required for running the annealed Langevin dynamics to estimate the normalizing constant within $\varepsilon$ relative error with high probability. The final bound depends on the action of the curve, specifically the integral of the squared metric derivative in Wasserstein-2 distance.
  • Building on the insights from the analysis of the continuous dynamics, in Sec. 5 we establish the first non-asymptotic oracle complexity bound for AIS, representing the first analysis of normalizing constant estimation algorithms without assuming a log-concave target distribution.
  • Finally, in Sec. 6, we point out a potential limitation of the geometric interpolation commonly used in annealing. To address this issue, we propose a novel algorithm based on reverse diffusion samplers and build up a framework for analyzing its oracle complexity.
Related works.
We briefly review some related works, and defer detailed discussion to App. H.
@@
  • Methods for normalizing constant estimation. We mainly discuss two classes of methods here. First, the equilibrium methods, such as TI (Kirkwood, 1935) and its variants (Brosse et al., 2018, Ge et al., 2020, Chehab et al., 2023, Kook and Vempala, 2024), which involve sampling sequentially from a series of equilibrium Markov transition kernels. Second, the non-equilibrium methods, such as AIS (Neal, 2001), which samples from a non-equilibrium SDE that gradually evolves from a prior distribution to the target distributions. In Sec. H.1, we show that TI is a special case of AIS using the “perfect” transition kernels.4 Recent years have also witnessed the emergence of learning-based non-equilibrium methods for normalizing constant estimation, which are typically byproducts of sampling algorithms (Zhang and Chen, 2022, Nüsken and Richter, 2021, Richter and Berner, 2024, Sun et al., 2024, Vargas et al., 2024, Albergo and Vanden-Eijnden, 2024, Blessing et al., 2025, Chen et al., 2025). Additionally, there are also several methods based on particle filtering (e.g., Kostov and Whiteley (2017), Jasra et al. (2018), Ruzayqat et al. (2022)).
@@
  • Variance reduction in JE and AIS. Our poof methodology focuses on the discrepancy between the sampling path measure and the reference path measure, which is related to the variance reduction technique in applying JE and AIS. For example, Vaikuntanathan and Jarzynski (2008) introduced the idea of escorted simulation, Hartmann et al. (2017) proposed a method for learning the optimal control protocol in JE through the variational characterization of free energy, and Doucet et al. (2022) leveraged score-based generative model to learn the optimal backward kernel. Quantifying the discrepancy between path measures is the core of our analysis.
@@
  • Complexity analysis for normalizing constant estimation. Chehab et al. (2023) studied the asymptotic statistical efficiency of the curve for TI measured by the asymptotic mean-squared error, and highlighted the advantage of the geometric interpolation. In terms of non-asymptotic analysis, existing works mainly rely on the isoperimetry of the target distribution. For instance, Andrieu et al. (2016) derived bounds of bias and variance for TI under Poincaré inequality, Brosse et al. (2018) provided complexity guarantees for TI under both strong and weak log-concavity conditions, while Ge et al. (2020) improved the complexity under strong log-concavity using multilevel Monte Carlo.
2 Preliminaries
Notations and definitions.
For $a,b\in\mathbb{R}$, let $\left[\!\left[a,b\right]\!\right]:=[a,b]\cap\mathbb{Z}$, $a\wedge b:=\min(a,b)$, and $a\vee b:=\max(a,b)$. For $a,b>0$, the notations $a\lesssim b$, $b\gtrsim a$, $a=O(b)$, $b=\Omega(a)$ indicate that $a\leq Cb$ for some constant $C>0$, and the notations $a\asymp b$, $a=\Theta(b)$ stand for $a\lesssim b\lesssim a$. $\widetilde{O}\left(\cdot\right),\widetilde{\Theta}\left(\cdot\right)$ hide logarithmic dependence in $O(\cdot),\Theta(\cdot)$. A function $U\in C^{2}(\mathbb{R}^{d})$ is $\alpha(>0)$-strongly-convex if $\nabla^{2}U\succeq\alpha I$, and is $\beta(>0)$-smooth if $-\beta I\preceq\nabla^{2}U\preceq\beta I$. We do not distinguish probability measures on $\mathbb{R}^{d}$ from their Lebesgue densities. For two probability measures $\mu,\nu$, the total-variation (TV) distance is $\operatorname{TV}(\mu,\nu)=\sup_{\textrm{measurable}~{}A}|\mu(A)-\nu(A)|$, and the Kullback-Leibler (KL) divergence is $\operatorname{KL}(\mu\|\nu)=\int\log\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\mathrm{d}\mu$. We call $\operatorname{\mathbb{E}}_{\mu}\|\cdot\|^{2}$ the second-order moment of $\mu$. Finally, a function $T:\mathbb{R}^{d}\times\mathbb{R}^{d}\to[0,{+\infty})$ is a transition kernel if for any $x$, $T(x,\cdot)$ is a p.d.f.
2.1 Stochastic Differential Equations and Girsanov Theorem
Throughout this paper, $(B_{t})$ and $(W_{t})$ represent standard Brownian motions (BM) on $\mathbb{R}^{d}$. For a stochastic differential equation (SDE) $X=(X_{t})_{t\in[0,T]}$ defined on $\Omega=C([0,T];\mathbb{R}^{d})$, the distribution of $X$ over $\Omega$ is called the path measure of $X$, defined by $\mathbb{P}^{X}$: measurable $A\subset\Omega\mapsto\operatorname{Pr}(X\in A)$. The following lemma , as a corollary of the Girsanov theorem (Üstünel and Zakai, 2013, Prop. 2.3.1 & Cor. 2.3.1), provides a method for computing the Radon-Nikodým (RN) derivative and KL divergence between two path measures, which serves as a key technical tool in our proof.
Lemma 1.
Assume we have the following two SDEs with $t\in[0,T]$:
We now define the backward SDE, which can be perceived as the time-reversal of a forward SDE. Given a BM $(B_{t})_{t\in[0,T]}$, let its time-reversal be $(B^{\leftarrow}_{t}:=B_{T-t})_{t\in[0,T]}$. We say that a process $(X^{\leftarrow}_{t})_{t\in[0,T]}$ satisfies the backward SDE
The forward and backward SDEs are related through the Nelson’s relation (Lem. 12), which also allows us to calculate the RN derivative between path measures of forward and backward SDEs (Lem. 15). We postpone the detailed derivations to App. A.
2.2 Wasserstein Distance, Metric Derivative, and Action
For two probability measures $\mu,\nu$ on $\mathbb{R}^{d}$ with finite second-order moments, the Wasserstein-2 ($\textbf{W}_{\textbf{2}}$) distance between $\mu$ and $\nu$ is defined as $W_{2}(\mu,\nu)=\inf_{\gamma\in\Pi(\mu,\nu)}\left(\int\|x-y\|^{2}\gamma(\mathrm{d}x,\mathrm{d}y)\right)^{\frac{1}{2}}$, where $\Pi(\mu,\nu)$ is the set of all couplings of $(\mu,\nu)$. The Brenier’s theorem states that when $\mu$ has a Lebesgue density, then there exists a unique coupling $\left(\operatorname{id}\times T_{\mu\to\nu}\right)_{\sharp}\mu$ that reaches the infimum. Here, $\sharp$ stands for the push-forward of a measure ($T_{\sharp}\mu(\cdot)=\mu(\{\omega:T(\omega)\in\cdot\})$), and $T_{\mu\to\nu}$ is known as the OT map from $\mu$ to $\nu$ and can be written as the gradient of a convex function.
Lemma 2.
Finally, we define the action of an AC curve of probability measures $(\rho_{t})_{t\in[a,b]}$ as $\int_{a}^{b}|\dot{\rho}|_{t}^{2}\mathrm{d}t$, which plays a key role in characterizing the efficiency of a curve for normalizing constant estimation. For basic properties of the action and its relation to isoperimetric inequalities such as log-Sobolev and Poincaré inequalities, we refer the reader to Guo et al. (2025, Lem. 3 & Ex. 1).
2.3 Langevin Diffusion and Langevin Monte Carlo
The (overdamped) Langevin diffusion (LD) with target distribution $\pi\propto\mathrm{e}^{-V}$ is the solution to
$\mathrm{d}X_{t}=-\nabla V(X_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},~{}t\in[0,\infty).$ (1)
Under mild regularity conditions, $\pi$ is the unique stationary distribution of this SDE, and when $\pi$ has good properties such as strong log-concavity, $X_{t}$ converges to $\pi$ in probability rapidly. In practice, when the closed-form solution of this SDE is unavailable, one usually leverages the Euler-Maruyama scheme to discretize Eq. 1, leading to the (overdamped) Langevin Monte Carlo (LMC) algorithm: with step size $h>0$, iterate the following update rule for $k=0,1,...$:
$X_{(k+1)h}=X_{kh}-h\nabla V(X_{kh})+\sqrt{2}(B_{(k+1)h}-B_{kh}),~{}\text{where}~{}B_{(k+1)h}-B_{kh}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\operatorname{\mathcal{N}}\left(0,hI\right).$ (2)
2.4 Reverse Diffusion Samplers
Inspired by score-based generative models (Song et al., 2021), recent advancements have led to the development of multimodal samplers based on reversing the Ornstein-Uhlenbeck (OU) process (Huang et al., 2024a, b, He et al., 2024, Vacher et al., 2025). In this paper, we collectively refer to these methods as the reverse diffusion samplers (RDS).
The following OU process transforms any target distribution $\pi$ into $\phi:=\operatorname{\mathcal{N}}\left(0,I\right)$ as $T\to\infty$:
$\mathrm{d}Y_{t}=-Y_{t}\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},~{}t\in[0,T];~{}Y_{0}\sim\pi,$ (3)
We denote the law of $Y_{t}$ by $\overline{\pi}_{t}$. The time-reversal $(Y^{\leftarrow}_{t}:=Y_{T-t}\sim\overline{\pi}_{T-t})_{t\in[0,T]}$ satisfies the SDE
Hence, to draw samples from $\pi$, it suffices to approximate the scores $\nabla\log\overline{\pi}_{t}$ and discretize Eq. 4, which can be implemented in various ways. For example, by Tweedie’s formula (Robbins, 1992), $\nabla\log\overline{\pi}_{t}$ is an affine function of $\operatorname{\mathbb{E}}(Y_{0}|Y_{t}=\cdot)$ (Eq. 34), while the law of $Y_{0}|Y_{t}=\cdot$ is analytically tractable (Eq. 35) and provably easier to sample from than the target $\pi$ (Huang et al., 2024a).
3 Problem Setting
Motivated by Brosse et al. (2018), Ge et al. (2020), given an accuracy threshold $\varepsilon\ll 1$, our goal is to study the complexity (measured by the number of calls to the oracles $V$ and $\nabla V$) required to obtain an estimator $\widehat{Z}$ of $Z$ such that with $\Omega(1)$ probability, the relative error is within $\varepsilon$:

### v2

@@ Abstract
Given an unnormalized probability density $\pi\propto\mathrm{e}^{-V}$, estimating its normalizing constant $Z=\int_{\mathbb{R}^{d}}\mathrm{e}^{-V(x)}\mathrm{d}x$ or free energy $F=-\log Z$ is a crucial problem in Bayesian statistics, statistical mechanics, and machine learning. It is challenging especially in high dimensions or when $\pi$ is multimodal. To mitigate the high variance of conventional importance sampling estimators, annealing-based methods such as Jarzynski equality and annealed importance sampling are commonly adopted, yet their quantitative complexity guarantees remain largely unexplored. We take a first step toward a non-asymptotic analysis of annealed importance sampling. In particular, we derive an oracle complexity of $\widetilde{O}\left(\frac{d\beta^{2}{\mathcal{A}}^{2}}{\varepsilon^{4}}\right)$ for estimating $Z$ within $\varepsilon$ relative error with high probability, where $\beta$ is the smoothness of $V$ and $\mathcal{A}$ denotes the action of a curve of probability measures interpolating $\pi$ and a tractable reference distribution. Our analysis, leveraging Girsanov theorem and optimal transport, does not explicitly require isoperimetric assumptions on the target distribution. Finally, to tackle the large action of the widely used geometric interpolation, we propose a new algorithm based on reverse diffusion samplers, establish a framework for analyzing its complexity, and empirically demonstrate its efficiency in tackling multimodality.
@@ 1 Introduction
We study the problem of estimating the normalizing constant $Z=\int_{\mathbb{R}^{d}}\widehat{\pi}(x)\mathrm{d}x$ of an unnormalized probability density function (p.d.f.) $\pi\propto\widehat{\pi}:=\mathrm{e}^{-V}$ on $\mathbb{R}^{d}$, so that $\pi(x)=\frac{1}{Z}\widehat{\pi}(x)$. The normalizing constant appears in various fields: in Bayesian statistics, when $\widehat{\pi}$ is the product of likelihood and prior, $Z$ is also referred to as the marginal likelihood or evidence (Gelman et al., 2013); in statistical mechanics, when $V$ is the Hamiltonian, [^1] Up to a multiplicative constant $\beta=\frac{1}{k_{\mathrm{B}}T}$ known as the thermodynamic beta, where $k_{\mathrm{B}}$ is the Boltzmann constant and $T$ is the temperature. When borrowing physical terminologies, we ignore this for simplicity. $Z$ is known as the partition function, and $F:=-\log Z$ is called the free energy (Chipot and Pohorille, 2007, Lelièvre et al., 2010, Pohorille et al., 2010). The task of normalizing constant estimation has numerous applications, including computing log-likelihoods in probabilistic models (Sohl-Dickstein and Culpepper, 2012), estimating free energy differences (Lelièvre et al., 2010), and training energy-based models in generative modeling (Song and Kingma, 2021, Carbone et al., 2023, Sander et al., 2025).
@@
Estimating normalizing constants is challenging in high dimensions or when $\pi$ is multimodal (i.e., $V$ has a complex landscape). Conventional approaches based on importance sampling (Meng and Wong, 1996) are widely adopted to tackle this problem, but they suffer from high variance due to the mismatch between the proposal and the target when $\pi$ is complicated (Chatterjee and Diaconis, 2018). To alleviate this issue, the technique of annealing tries constructing a sequence of intermediate distributions that bridge these two distributions, which motivates several popular methods including path sampling (Chen and Shao, 1997, Gelman and Meng, 1998), annealed importance sampling (AIS, Neal (2001)), and sequential Monte Carlo (SMC, Doucet et al. (2000), Del Moral et al. (2006), Syed et al. (2024)) in statistics literature, as well as thermodynamic integration (TI, Kirkwood (1935)) and Jarzynski equality (JE, Jarzynski (1997), Ge and Jiang (2008), Hartmann et al. (2019)) in statistical mechanics literature. In particular, JE points out the connection between the free energy difference between two states and the work done over a series of trajectories linking these two states, while AIS constructs a sequence of intermediate distributions and estimates the normalizing constant by importance sampling over these distributions. These two methods are our primary focus in this paper.
@@
Despite the empirical success of annealing-based methods (Ma et al., 2013, Krause et al., 2020, Mazzanti and Romero, 2020, Yasuda and Takahashi, 2022, Chen and Ying, 2024, Schönle et al., 2025), the theoretical understanding of their performance is still limited. Existing works for importance sampling mainly focus on the asymptotic bias and variance of the estimator (Meng and Wong, 1996, Gelman and Meng, 1998), while works on JE usually simplify the problem by assuming the work follows simple distributions (e.g., Gaussian or gamma) (Echeverria and Amzel, 2012, Arrar et al., 2019). Moreover, only analyses asymptotic in the number of particles derived from central limit theorem exist (Lelièvre et al., 2010, Sec. 4.1). This paper aims to establish a rigorous non-asymptotic analysis of estimators based on JE and AIS, while introducing minimal assumptions on the target distribution. We also propose a new algorithm based on reverse diffusion samplers to tackle a shortcoming of AIS.
Contributions. 1. We discover a novel strategy for analyzing the complexity of normalizing constant estimation, applicable to a wide range of target distributions (Assumps.˜2 and 1) that may not satisfy isoperimetric conditions such as log-concavity. 2. In Sec.˜3, we study JE and prove an upper bound on the time required for running the annealed Langevin dynamics to estimate the normalizing constant within $\varepsilon$ relative error with high probability. The final bound depends on the action (the integral of the squared metric derivative in Wasserstein-2 distance) of the curve. 3. Building on the insights from this analysis of the continuous dynamics, in Sec.˜4 we establish the first non-asymptotic oracle complexity bound for AIS, representing the first analysis of normalizing constant estimation algorithms without assuming a log-concave target distribution. 4. Finally, in Sec.˜5, we first point out a potential limitation of the commonly used geometric interpolation, which provides a quantitative explanation of the mass teleportation phenomenon. We then propose a series of new algorithms based on reverse diffusion samplers and formalize a framework for analyzing its oracle complexity. Our experimental results demonstrate the superiority of the proposed algorithm over AIS in overcoming multimodality.
Related Works. Below, we summarize the related works in four aspects.
@@
I. Methods for normalizing constant estimation. We mainly discuss two classes of methods here. First, the equilibrium methods, such as TI (Kirkwood, 1935) and its variants (Brosse et al., 2018, Ge et al., 2020, Chehab et al., 2023, Kook and Vempala, 2025), which involve sampling sequentially from a series of equilibrium Markov transition kernels. Second, the non-equilibrium methods, such as AIS (Neal, 2001), which samples from a non-equilibrium SDE that gradually evolves from a prior distribution to the target distributions. In Sec.˜H.1, we show that TI is a special case of AIS using the “perfect” transition kernels. Recent years have also witnessed the emergence of learning-based non-equilibrium methods, which are typically byproducts of neural samplers (Nüsken and Richter, 2021, Zhang and Chen, 2022, Máté and Fleuret, 2023, Richter and Berner, 2024, Sun et al., 2024, Vargas et al., 2024, Máté et al., 2024, Albergo and Vanden-Eijnden, 2025, Blessing et al., 2025, Chen et al., 2025, He et al., 2025). Finally, there are also methods based on particle filtering (Kostov and Whiteley, 2017, Jasra et al., 2018, Ruzayqat et al., 2022).
@@
II. Variance reduction in JE and AIS. Our proof methodology focuses on the discrepancy between the sampling path measure and the reference path measure, which is related to the variance reduction technique in applying JE and AIS. For example, Vaikuntanathan and Jarzynski (2008) introduced the idea of escorted simulation, Hartmann et al. (2017) proposed a method for learning the optimal control protocol in JE through the variational characterization of free energy, and Doucet et al. (2022) leveraged score-based generative model to learn the optimal backward kernel. Quantifying the discrepancy between path measures is the core of our analysis.
@@
III. Complexity analysis for normalizing constant estimation. Chehab et al. (2023) studied the asymptotic statistical efficiency of the curve for TI measured by the asymptotic mean-squared error, and highlighted the advantage of the geometric interpolation. In terms of non-asymptotic analysis, existing works mainly rely on the isoperimetry of the target distribution. For instance, Andrieu et al. (2016) derived bounds of bias and variance for TI under Poincaré inequality (PI), Brosse et al. (2018) provided complexity guarantees for TI under both strong and weak log-concavity conditions, while Ge et al. (2020) improved the complexity under strong log-concavity using multilevel Monte Carlo.
IV. Complexity analysis of sampling beyond isoperimetry. Our analysis of estimating normalizing constants of non-log-concave distributions is also closely related to the study of sampling beyond log-concavity. In general, such problems are NP hard (Ge et al., 2018, He and Zhang, 2025). Existing works providing convergence guarantees have leveraged more general isoperimetric inequalities such as weak PI (Mousavi-Hosseini et al., 2023), tried to establish convergence in weaker notions (Balasubramanian et al., 2022, Cheng et al., 2023), or utilized denoising diffusion model (Huang et al., 2024a, He et al., 2024). We highlight Guo et al. (2025) that this paper mainly draws inspiration from, which introduced the action of a curve in quantifying the convergence of annealed sampling. While they focused on sampling and presented cases where annealing works, we extend the analysis to a conceptually different task, and further establish lower bounds on the action of the commonly used geometric interpolation, motivating a new algorithm based on reverse diffusion samplers.
Notations and Preliminaries. For brevity, we integrate the required background information into the main text, with a detailed exposition available in App.˜A. The notations can also be found there.
2 Preliminaries and Problem Setting
To motivate the study of normalizing constant estimation, we first present several examples.
Example 1. [Free energy difference.] In many statistical physics problems (Lelièvre et al., 2010), given two energy functions $U_{0},U_{1}$ (possibly linked through some thermodynamic process), one is often interested in estimating the free energy difference $\Delta F:=-\frac{1}{\beta}\log(\int\mathrm{e}^{-\beta U_{1}}\mathrm{d}x/\int\mathrm{e}^{-\beta U_{0}}\mathrm{d}x)$, which is related with the normalizing constant of the distributions $\pi_{i}\propto\mathrm{e}^{-\beta U_{i}}$.
Example 2. [Likelihood in latent variable models.] In latent variable models such as variational autoencoders (Kingma and Welling, 2013), a common evaluation metric is the marginal likelihood of a data point $x$, $p_{\theta}(x)=\int p_{\theta}(x|z)p(z)\mathrm{d}z$. This is nothing but the normalizing constant of the posterior distribution of the latent variable $z$ given data $x$, $p_{\theta}(z|x)\propto_{z}p_{\theta}(x|z)p(z)$.
Example 3. [Volume of convex bodies.] In theoretical computer science, a classical problem is to estimate the volume of a convex body $\mathcal{K}$ (Dyer et al., 1991, Cousins and Vempala, 2018, Kook et al., 2024), which is equivalent to the normalizing constant of the uniform distribution on $\mathcal{K}$, $\pi\propto 1_{\mathcal{K}}$.
Building on prior theoretical results (Brosse et al., 2018, Ge et al., 2020), we study the oracle complexity of estimating normalizing constant of a density under the following criterion:
Aim: Given a density $\pi\propto\widehat{\pi}:=\mathrm{e}^{-V}$ on $\mathbb{R}^{d}$, bound the complexity of obtaining an estimator $\widehat{Z}$ of $Z=\int_{\mathbb{R}^{d}}\widehat{\pi}(x)\mathrm{d}x$ such that with constant probability, the relative error is within $\varepsilon(\ll 1)$: